when news makes individual (less/more) patient about discounting future cash flows, this can lead to an in (increse/decrese) asset price?

Please choose answer from two brackets.
Answer #1

When news makes individual more patient about discounting future cash flows , this can lead to an decrease in asset price

Reason: When investor become more patient , it means that they expect cash flow to occur in latter year than now and thus price of asset decreases. For example say a earlier investor were expecting $ 100 cash flow in year 1 and now they are expecting cash flow in year 2

Assuming rate to be 10%

Value of asset when cash flow was expected to occur in year 1 = 100/1.1 = 90.91 $

Value of asset when cash flow was expected to occur in year 2 = 100/1.1^2 = 100/1.21 = 82.64 $